OpenStudy (anonymous):

Multiply the following: 7x^2√2x^8 * 6x^4√23x^9 * 3x^16

OpenStudy (anonymous):

564159456 sqrt(23) x^(61/2)

OpenStudy (anonymous):

i do not have that

OpenStudy (anonymous):

\[7x^2 * \sqrt{2} x^4*6x^4*\sqrt{23}*x^{9/2}*3x^{16}\]Remember that when exponents of the same base are multiplied, they exponents add. The square root has an exponent of 1/2.

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

(54 x^22) 2 x^7 = 108 x^29

OpenStudy (anonymous):

Therefore, we would have \[(7*\sqrt{2}*6*\sqrt{24} * 3)*x^{2+4+(9/2) + 16}\]

OpenStudy (anonymous):

i have like 126x^30

OpenStudy (anonymous):

kayser, let me verify that equation you typed. Is it \[7x^2*\sqrt{2x^8}*6x^4*\sqrt{23x^9}*3x^{16}\]

OpenStudy (anonymous):

sir the 6x^4 is with √23x^9

OpenStudy (anonymous):

it is like this 6x^4√23x^9

OpenStudy (anonymous):

\[6x^{4\sqrt{23x^9}}\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what would it be then

OpenStudy (anonymous):

Give me a second to type it up.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[7x^{2*\sqrt{2}*x^4}*6x^{4\sqrt{23} \sqrt{x^9}}*3x^{16}\]This further simplifies to \[(7*6*3)^{(2*\sqrt{2}*x^4 + 4*\sqrt{23}x^{9/2}+16)}\] I don't think we can go any further in the exponent.

OpenStudy (anonymous):

kayser, please see my updated solution to the inequality question you asked earlier.